# World In Your Hands: A Large-Scale and Open-source Ecosystem for Learning Human-centric Manipulation in the Wild

### v1

@@ 3.1 Human-centric Vision-Language (HVL)
In this section, we construct three evaluation tasks based on the vision-language annotations introduced in Section 2. These tasks are designed to evaluate whether existing Multimodal Large Language Models (MLLMs) can achieve a genuine understanding of human-centric manipulation tasks in highly generalized in-the-wild environments.
@@
We define the following three evaluation tasks and test several closed-source and open-source MLLMs, including GPT-4o [hurst2024gpt], Doubao-seed [bai2024seed], and Qwen [yang2025qwen3]. Detailed experimental setups and data statistics are provided in the supplementary material. (1) Spatial Referring (SR): Given a human manipulation image and a description of the spatial relationships among multiple objects in the scene, the model is required to output the pixel coordinates of a specified region. Performance is measured by the probability that the predicted point lies within the target area. (2) Subtask Prediction (SP): Given a piece of human manipulation video, the current task description, and four candidate options for the next action, the model must select the correct subsequent action. (3) Completion Verification (CV): Given a piece of human manipulation video and the current task description, the model needs to determine whether the task has been completed.
As shown in Table 2, Qwen3-VL-Plus, which was pre-trained on spatial understanding data, achieves the best performance on the Spatial Referring task. However, scores of all MLLMs are below 0.5, indicating that current MLLMs still lack sufficient capability in interpreting fine-grained spatial relationships in manipulation scenarios. On the Completion Verification task, all models perform slightly above 0.5, suggesting that their ability to accurately judge task completion in such contexts remains limited. These experimental results collectively demonstrate that general-purpose, vanilla VLLMs still fall short in embodied tasks. Our WiYH dataset, with its rich annotations of human activities and spatial relationships, presents a promising foundation for future embodied pre-training.
| Model | SR (%) | SP (%) | CV (%) |
@@
3.2 Human-centric World Model (HWM)
| CogVideo [hong2022cogvideo] | DynamiCrafter [xing2024dynamicrafter] | |||
| Metric | w/o WiYH | w/ WiYH | w/o WiYH | w/ WiYH |
| Consistency | 88.2 (+7.6) | 91.5 (+6.2) | ||
| Smoothness | 98.4 (+4.9) | 98.6 (+1.2) | ||
| Dynamic | 78.5 (+15.6) | 89.6 (+14.8) | ||
| Quality | 67.6 (+7.5) | 70.6 (+7.1) | ||
3.2.1 Language-conditioned Video Generation
This section evaluates the contribution of our proposed WiYH dataset to enhancing video generation performance in human-centric manipulation tasks. We fine-tuned two baseline models, CogVideo [hong2022cogvideo] and DynamiCrafter [xing2024dynamicrafter], initialized with their original weights, using WiYH under the image-and-text-to-video setting. During fine-tuning, all input videos were resized to 480p resolution.
Quantitative results are summarized in Table 3. We use four metrics from VBench [huang2024vbench]—consistency, smoothness, dynamic (motion intensity), and quality—to assess motion plausibility, temporal coherence, and visual fidelity. After fine-tuning with the WiYH dataset, both baseline models show clear improvements across all metrics, with the most significant gains observed in motion intensity and temporal consistency.
3.2.2 Gaussian Splatting 4D Reconstruction
@@
Figure 6 presents the 4D reconstruction results at various time steps, showcasing the dataset’s potential for dynamic scene modeling. By capturing high-quality dynamic scenes, WiYH can help leverage Gaussian Splatting (GS) as an effective 3D tokenizer, advance spatial understanding, and enrich Real2Sim pipelines.
@@
3.3 Human-centric VLA Training (HVT)
In this section, we evaluate several Vision-Language-Action (VLA) algorithms trained on the WiYH dataset to explore whether its visual-action-language tri-modal support enables effective training of vision-language-action-aligned manipulation policies.
Specifically, we use GR00T [bjorck2025gr00t] and OpenVLA-OFT [kim2025fine] as baseline algorithms, training each to predict the future 1.6-second motion trajectory (16 points at 10 Hz) of the human wrist from the dataset. We sample 10% of the data from each scenario for training and assemble a test set of 400 non-overlapping episodes. Each baseline model is trained for 30,000 steps with a total batch size of 128 on 8 NVIDIA A100 GPUs.
As shown in Table 5, training with human-centric data consistently reduces the error in tracking the ground-truth trajectories on the test set across both mainstream VLA frameworks. These results demonstrate WiYH’s effectiveness for training VLA algorithms. We believe the WiYH dataset can serve as a foundational resource for validating scaling laws in human-centric VLA pre-training.
| Model | Mean Error (cm) | ||
| 0 Steps | 15k Steps | 30k Steps | |
| OpenVLA-OFT [kim2025fine] | 4.75 | 1.28 | 1.12 |
| GR00T [bjorck2025gr00t] | 4.54 | 1.31 | 1.07 |
4 Cross-Embodiment Analysis
In this section, we conduct co-training experiments using cross-embodiment data to investigate the effectiveness of the human video data.
4.1 Implementation Details
We set up two scenarios: a single-object scene and a multi-object cluttered scene. In this experiment, we employ a co-training policy using data collected by an improved version of DexUMI [xu2025dexumi] and from Human Video demonstrations. Figure 8 shows representative scene examples and the data collection setup. For UMI, data collection always started from a limited number of initial poses. The improved UMI system is capable of capturing trajectories with millimeter-level accuracy and was exclusively used to collect high-precision manipulation data in the single-object scene, resulting in fine-grained motion trajectories. In contrast, Human Video contains only manipulation sequences from multi-object cluttered scenes, providing data that spans a larger action space, exhibits a broader distribution of motions, and covers a greater diversity of object observations.
@@ 4.2 Results and In-depth Discussion
In the single-object scenario, the robot arm begins execution from a random initial pose during testing, which differs from the initial pose used in the UMI dataset. In the multi-object cluttered scenario, we primarily evaluate the performance of the co-training strategy under conditions of varying scene height, changes in the target object’s pose, and increased occlusion. The configurations of these test scenarios and qualitative experimental results are illustrated in Figure 7.
Quantitative results are summarized in Table 6. (1) In the single-object scenario, co-training the policy with both UMI data and Human Video data leads to an improvement in the success rate of more than 13%. Although the motion precision in the human data is limited, its diverse action distribution enhances the policy’s generalization capability to unseen initial states. (2) Policies trained solely on UMI data exhibit almost no generalization ability in complex scenarios; however, after co-training with human data, the SR increases by 52%. This demonstrates that incorporating human data significantly improves the policy’s ability to interpret complex scenes. Therefore, the co-training approach effectively introduces both the action domain and observation domain of human data into the robot dataset, thereby enhancing the generalization capability of the learned policy.
| Scenario | Data Sources and Amounts | SR (%) | |
| UMI (clips) | Human Video (clips) | ||
@@ 5 Conclusion
This paper presents the WiYH ecosystem, a cornerstone for embodied intelligence research in real-world environments. We make three key contributions: the Oracle Suite, a wearable system enabling in-the-wild data capture with integrated markerless auto-labeling; the large-scale WiYH Dataset, offering diverse, skilled human manipulation sequences with rich multi-modal streams; and a suite of Comprehensive Benchmarks, supported by extensive atomic-action and vision-language annotations, for holistic evaluation from perception to physical interaction. Our benchmarks reveal that current models still lack the fine-grained spatial and causal reasoning required for robust embodied operation. The WiYH dataset, with its unique scale, and rich annotations, is designed to bridge this gap. We believe this work provides an essential foundation for developing more capable and generalizable embodied agents.
Appendix A World in Your Hand (WiYH) Dataset
A.1 Detailed Task-Skill Counting Statistics
We further supplement the visualization of the number of skills corresponding to each task. Due to space limitations, we report statistics based on only 25% of the dataset, as shown in Figure. 16. The full visualization analysis will be updated subsequently on the dataset website.
A.2 Additional annotation examples visualizations
We provide additional annotation visualizations for the remaining 6 scenes, as shown in Figure 15.
A.2.1 Data Annotation Details.
Raw videos are structured into a three-tier hierarchy: Scene - Task - Subtask by Annotators. Each subtask corresponds to the minimal object-centered action unit or a temporally cohesive action sequence. (1) Subtask Annotation:
Actions: Drawn from predefined operation sets (e.g., PICK, PLACE_INTO, SEARCH).
Target Object: Semantic structure describing action objects and positional relationships (e.g., [object, container]).
@@ A.2 Additional annotation examples visualizations
(2) Subtask Prediction chain-of-thought (CoT): With the subtask annotation, we select tasks of moderate length (containing 4-10 subtasks). We annotate them with Chain-of-Thought data, providing reasoning that decomposes the task into several logically connected subtasks. The annotation pipeline for CoT is illustrated in Figure 10. Initially, we provide the model (Qwen2.5-VL) with both the current subtask information (video and text) and the ground-truth text of the next subtask. The model is then prompted to perform posterior attribution, explaining why the specified next subtask logically follows. Subsequently, the attribution is appended to the input and presented again to Qwen2.5-VL, which synthesizes it into a structured causal reasoning process. Finally, the CoT segment is extracted and used to prompt Qwen-3, which is instructed to predict an answer. This predicted answer is matched against the ground-truth answer to validate the completeness and soundness of the CoT labeling.
Appendix B Benchmark Details
B.1 Human-centric Vision-Language
B.1.1 Benchmark Task Descriptions.
The Human-centric Vision-Language Benchmark in Section 3.1 is constructed by sampling raw data from WiYH, followed by manual annotation, including Spatial Referring, Subtask Prediction, and Completion Verification, with distributions illustrated in Figure 9. (1) The Spatial Referring is designed as a spatial region grounding form: given relevant image frames, annotators delineate ROI areas using polygons according to the given questions. It evaluates spatial understanding by prompting the model to output a set of points representing the spatial region referenced in the question. The primary metric is the proportion of these points that fall within the ground-truth polygon. (2) The Subtask Prediction adopts a multiple-choice question format. Using Qwen2.5-VL-72B, we first generate 9 negative candidate options based on video segments and ground truth, spanning action, object, and orientation dimensions. Annotators then manually select the top three hard negative samples, which, together with the correct answer, form the final candidate set. (3) The Completion Verification is designed as binary questions: given a video segment randomly truncated from the end, annotators determine whether the subtask has been completed. And this evaluation metric is the selection accuracy rate.
B.1.2 Benchmark Tasks Visualization.
Examples of the three tasks are shown in Figure 11.
B.2 Human-centric World Model
B.2.1 Language-conditioned Video Generation
Baseline. We validate the effectiveness of our WiYH dataset using video generation baselines with different architectures, including transformer (CogVideo [hong2022cogvideo]) and U-Net (DynamiCrafter [xing2024dynamicrafter]).
Evaluation Metrics. Following VBench [huang2024vbench], the evaluation metrics are as follows:
Implementation Details. During fine-tuning, all input videos are resized to 480p resolution. For CogVideo, we use the pre-trained image-to-video version with a temporal length of 49, a learning rate of 2e-5, and a batch size of 16. For Dynamicrafter, we leverage the model pre-trained at 512 resolution, setting the temporal length to 48, the learning rate to 1e-5, and the batch size to 16.
B.2.2 Gaussian Splatting 4D Reconstruction
@@
Implementation Details. During optimization, all input videos are processed at a resolution of 960×720. We represent each scene with 40k dynamic Gaussians and 100k static Gaussians and use B = 10 shared SE(3) motion bases for all points. We optimize the model with Adam, running 1k iterations for the initialization stage and 500 training epochs, with learning rates of for Gaussian means and SE(3) bases, for scales, for rotations, and for colors, opacities, and motion coefficients. Training on a 300-frame sequence takes about 2 hours on a single A100 GPU and supports real-time rendering at around 140 FPS.
B.3 Details of Cross-Embodiment Analysis.
In this section, we present the details of our cross-embodiment experiments in Section 4, Cross-Embodiment Analysis, of the main text. As shown in Figure 14, we demonstrate 11 task variants under two scenarios: Single-Object Scene and Multi-Object Cluttered Scene. Figures 12 and 13 illustrate the success rates and failure causes for each task in these two scenarios, respectively.
As shown in Figures 12 and 13, on most tasks, the incorporation of human data leads to a general increase in success rate. In the Single-Object Scene, the diverse action distribution of human data enhanced the policy’s generalization capability to unseen initial states, enabling successful grasping at different positions or over long distances. In the Multi-Object Cluttered Scene, policies trained solely on UMI data demonstrated almost no generalization ability in complex scenarios. When factors such as object height or layout changed, policies trained only on UMI data failed, whereas the inclusion of human data gradually enabled success by effectively introducing both the action and observation domains of human data into the robot dataset.
Furthermore, the consistently low success rates across all methods in certain tasks highlight important directions for future research: how to further improve the collection precision of human data and how to design efficient algorithms that can better transfer knowledge from human data to dexterous manipulation tasks.
Appendix C Ethics and Privacy Statement

### v3

@@ 4.1 Benchmark: Human-centric Vision-Language (HVL)
In this section, we construct three evaluation tasks based on the vision-language annotations introduced in Section 3. These tasks are designed to evaluate whether existing Multimodal Large Language Models (MLLMs) can achieve a genuine understanding of human-centric manipulation tasks in highly generalized, in-the-wild environments.
@@
We define the following three evaluation tasks and test several closed-source and open-source MLLMs, including GPT-4o [hurst2024gpt], Doubao-seed [bai2024seed], and Qwen [yang2025qwen3]. Detailed experimental setups and data statistics are provided in the supplementary material. (1) Spatial Referring (S.R.): Given a human manipulation image and a description of the spatial relationships among multiple objects in the scene, the model is required to output the pixel coordinates of a specified region. Performance is measured by the probability that the predicted point lies within the target area. (2) Subtask Prediction (S.P.): Given a piece of human manipulation video, the current task description, and four candidate options for the next action, the model must select the correct subsequent action. (3) Completion Verification (C.V.): Given a piece of human manipulation video and the current task description, the model needs to determine whether the task has been completed. Evaluation details will be provided in the supplementary material. As shown in Table 2, scores for the S.R. task and C.V. task are generally low, whereas scores for the S.P. task remain relatively high. This discrepancy reflects the characteristics of existing VLMs, which demonstrate competent performance in high-level task decomposition but exhibit limitations in dynamic progress modeling and spatial understanding in specific embodied tasks. Specifically, on the S.R. task, scores of all VLMs are below 0.5, and on the C.V. task, all models perform slightly above 0.5, suggesting that their abilities in such contexts remain limited. Notably, Qwen3-VL-Plus, which was pre-trained on spatial-understanding data, achieves the best performance on the S.R. task. These experimental results collectively demonstrate that general-purpose, vanilla VLMs still fall short in embodied tasks. Our WIYH dataset, with its rich annotations of human activities, scene depth, and object masks, presents a promising foundation for future embodied pre-training.
| Model | S.R. (%) | S.P. (%) | C.V. (%) |
@@
4.2 Application: Human-centric World Modeling
@@
Figure 7 presents the 4D reconstruction results at various time steps, showcasing the dataset’s potential for dynamic scene modeling. By capturing high-quality dynamic scenes, WIYH can help leverage Gaussian Splatting (GS) as an effective 3D tokenizer, advance spatial understanding, and enrich Real2Sim pipelines and neural-network-based simulation.
In addition, we further present the experimental results of video generation in the supplementary materials.
@@
5 Human-centric Manipulation
5.1 Experiment Objective and Implementation
In this section, our objective is to validate the efficacy of the WIYH data collection system and dataset for robotics manipulation. To this end, we formulate two core research questions. (1) Does utilizing the WIYH dataset as pre-training data enhance the embodied manipulation capabilities of the VLA model? (2) Does retargeting WIYH data to dexterous hand actions improve dexterous manipulation performance? To address these questions, we design two real-world manipulation experiments: (1) cross-embodiment pretraining and (2) retargeting co-training.
In the cross-embodiment pretraining experiment, we adopt StarVLA-PI [starvla2025] as the base model. We pre-train StarVLA-PI using WIYH data with atomic instruction annotations for both VLM and VLA data. For post-training, we select two tasks in a real-world environment: rose insertion and gift packing, and we collect 150 gripper-based teleoperation manipulation samples for each task, as shown on the left side of Figure 8. We compare the success rates across three settings, including VLM pre-training, VLA pre-training, and without pre-training. VLA pretraining refers to loading all pre-trained VLA weights during post-training, while VLM pretraining denotes loading only the VLM weights. During evaluation, each setting of each task is tested 10 times with object positions and target objects varied in each trial.
In the retargeting co-training experiment, we construct a combined dataset using UMI robot data collected in constrained environments and data of the same task collected using the Oracle Suite. The robot system employs a low-degree-of-freedom dexterous hand as the end-effector , and the hand data from is retargeted to match the degrees of freedom of . We evaluate the effectiveness of the human-centric data collected by Oracle Suite across various dexterous grasping tasks. Implementation details will be provided in the supplementary materials. Notably, the right side of figure 8 shows scene examples and the data collection setup. The collected robot data always starts from a limited number of initial poses in the single-object scene. In contrast, human-centric data contains only manipulation sequences from multi-object cluttered scenes, making a larger observation domain and action space than .
@@ 5.2 Results and In-depth Discussion
In this section, we analyze the results of two experiments to demonstrate how WIYH empowers robotic manipulation.
For the cross-embodiment pretraining experiments, we demonstrate the quantitative result and the training loss in Table 4 and in Figure 9, respectively. As indicated by success rates, utilizing WIYH for VLA pre-training yields the most significant performance improvements, from 15% to 70%. Besides, as illustrated in Figure 9, VLA pre-training achieves lower loss and smoother gradient convergence than VLM pre-training and no pre-training. This efficacy remains robust given the cross-embodiment gap between the pre-training data (human hands) and post-training data (robot gripper).
| Method | Rose Insertion | Gift Packing | AVG |
| w/o Pretrain | 1/10 | 2/10 | 15% |
| VLM Pretrain | 3/10 | 3/10 | 30% |
| VLA Pretrain | 6/10 | 8/10 | 70% |
For the retargeting co-training experiments, we demonstrate the quantitative and qualitative results in Table 5 and in Figure 10, respectively. In the single-object scenario, the robot arm begins execution from a random initial pose during testing, which differs from the initial pose used in the robot data . In the multi-object cluttered scenario, we primarily evaluate the performance of the co-training strategy under conditions of varying scene height, changes in the target object’s pose, and increased occlusion. The configurations of these test scenarios and qualitative experimental results are illustrated in Figure 10. Quantitative results are summarized in Table 5. (1) In the single-object scenario, co-training the policy with both robot data and human-centric data leads to an improvement in the success rate of more than 13%. Although the motion precision in the human-centric data is limited, its diverse action distribution enhances the policy’s generalization capability to unseen initial states. (2) Policies trained solely on robot data exhibit minimal generalization capability in complex scenarios. Simply increasing data quantity from 200 clips to 500 clips yields marginal performance improvement from 0% to 8%. However, after co-training with human-centric data, the success rate increases by 52%. This demonstrates that incorporating human-centric data significantly improves the policy’s ability to interpret complex scenes. Therefore, the co-training approach effectively introduces both the action domain and observation domain of human-centric data into the robot dataset, thereby enhancing the generalization capability of the learned policy.
| Scenario ă | Data Sources and Amounts (Clips) ă | SR (%) | |
| ă | Robot data amount | Human data amount | ă |
@@ 6 Conclusion
In this paper, we present the WIYH ecosystem, a cornerstone for embodied intelligence research in real-world environments. We make three key contributions: the Oracle Suite, a wearable system enabling in-the-wild data capture with integrated markerless auto-labeling; the large-scale WIYH dataset, offering diverse, skilled human manipulation sequences with rich multimodal streams; and a suite of comprehensive benchmarks, supported by extensive atomic-action and vision-language annotations, for holistic evaluation from perception to physical interaction. Our benchmarks reveal that current models still lack the fine-grained spatial and causal reasoning required for robust embodied operation. The WIYH dataset, with its unique scale and rich annotations, is designed to bridge this gap. In the future, we will open-source the whole dataset and hardware design to promote human-centric manipulation research.
Appendix
Appendix 0.A World In Your Hands Ecosystem
0.A.1 Hardware Details
Camera: We primarily use two fisheye cameras below each wrist and two fisheye cameras on the chest as the raw RGB signal sources. These cameras feature a 180° field of view with an original resolution of 1536 × 1920 pixels.
Tactile: We utilize resistive pressure sensors with a resolution of 5 mN and a range of 0.2 N to 50 N distributed across the fingertips of all fingers.
Autolabel Pipeline: The autolabeling pipeline involves online and offline algorithms. Specifically, the online localization module performs lightweight coarse positioning using visual-inertial odometry (VIO) algorithms [qin2018robust] and IR detection, while the offline stage refines these coarse estimates through Structure-from-Motion (SfM) algorithms [schoenberger2016sfm].
0.A.2 Data Collection Efficiency
The Oracle Suite pipeline follows a Collection, Annotation, Inspection, Archiving workflow for scalable data production. A single collector operating 8 hours per day generates approximately 1.8 TB of valid multimodal data. At scale, 100 collectors over 30 days yield roughly 5.4 PB of raw data (stored in cloud repositories). Efficiency is further improved through standardized capture protocols and automated cloud-based annotation.
We compare the efficiency of different data collection paradigms in Table 6. Efficiency: taking the rose insertion task in our experiments as an example, given 8 hours as collection time, teleoperation can collect about 150 data clips; UMI-like methods, such as DexUMI [xu2025dexumi], can collect about 400 data clips; our oracle suite can collect 720 data clips; and VR can collect around 800 data clips. Quality: VR-based hand tracking only provides 2D hand skeletons and suffers from visual occlusion. In contrast. In comparison, the Oracle Suite provides accurate 3D hand skeletons, even if under visual occlusion.
| Teleop | DexUMI | VR | Oracle Suit (Ours) | |
| Efficiency | ||||
| Quality | ||||
| End-effector | Dexterous | Dexterous | Hands | Hands |
Appendix 0.B World In Your Hands Dataset
0.B.1 Detailed Task-Skill Counting Statistics
We further supplement the visualization of the number of skills corresponding to each task. Due to space limitations, we report statistics based on only 25% of the dataset, as shown in Figure 11. Please zoom in electronically to inspect the details. The full visualization analysis will be updated subsequently on the dataset website.
0.B.2 Data Annotation Details.
Raw videos are structured into a three-tier hierarchy: Scene - Task - Subtask by annotators. Each subtask corresponds to the minimal object-centered action unit or a temporally cohesive action sequence. (1) Atomic (Subtask) Annotation:
Actions (Skills): Drawn from predefined operation sets (e.g., PICK, PLACE INTO, SEARCH).
Target Object: Semantic structure describing objects (e.g., [cloth, container]).
@@ 0.B.1 Detailed Task-Skill Counting Statistics
(2) Chain-of-Thought (CoT) Reasoning: With the subtask annotation, we select tasks of moderate length (containing 4-10 subtasks). We annotate them with Chain-of-Thought data, providing reasoning that decomposes the task into several logically connected subtasks. The annotation pipeline for CoT is illustrated in Figure 12. Initially, we provide the model (Qwen2.5-VL-72B [Qwen2.5-VL]) with both the current subtask information (video and subtask instruction) and the ground-truth instruction of the next subtask. The model is then prompted to perform posterior attribution, explaining why the specified next subtask logically follows. Subsequently, the attribution is appended to the input and presented again to Qwen2.5-VL-72B, synthesizing it into a structured causal reasoning process. Next, the CoT segment is extracted and used to prompt Qwen3 [yang2025qwen3], which is instructed to predict an answer. This predicted answer is matched against the ground-truth answer to automatically validate the completeness and soundness of the CoT labeling. Finally, human review and correction of erroneous annotations are performed.
0.B.3 More visualization.
We present more visualizations about the annotations of our dataset in Figure 13. Annotations of the Human-centric Vision-Language tasks are shown in Figure 14.
Appendix 0.C Benchmark And Application Details
0.C.1 Human-centric Vision-Language
0.C.1.1 Benchmark Task Descriptions.
The Human-centric Vision-Language Benchmark in Section 4.1 is constructed by sampling raw data from WIYH, followed by manual annotation, including Spatial Referring, Subtask Prediction, and Completion Verification, with the amounts and distributions illustrated in Figure 15. In these tasks, all VLMs are directly tested without any training.
(1) The Spatial Referring is designed as a spatial region grounding form: given relevant image frames, annotators delineate ROI areas using polygons according to the given questions. Specifically, given the question as prompts, the VLM outputs a set of points representing the spatial region referenced in the question. We denote the set of points that fall within the ground-truth region as . We evaluate the VLM’s spatial understanding by computing the proportion of the points that fall within the ground-truth region.
where is the number of and is the number of . A is a scalar in the range , with values closer to 1 indicating stronger spatial understanding capability of the VLM. This task is designed to evaluate the capability of VLM models to understand spatial referring in human-centric scenarios.
(2) The Subtask Prediction adopts a multiple-choice question format. We use the Qwen2.5-VL-72B tool to first generate nine negative candidate options based on video clips and ground truth values, covering both action and target object dimensions. Then, the annotators manually select three reasonable negative samples that fit the current scenario. These samples, along with the correct answers, constitute the final candidate set. That is, each question has four options, and the accuracy of random selection was approximately 25%. This task is designed to evaluate the capability of VLM models to perform long-horizon task planning and decomposition in human-centric scenarios.
(3) The Completion Verification is designed as binary questions: given a video segment randomly truncated from the end, annotators determine whether the subtask has been completed. The evaluation metric is the selection accuracy rate. The accuracy of random selection was approximately 50%. This task is designed to evaluate the capability of VLM models to understand the dynamic process of manipulation tasks in human-centric scenarios.
0.C.1.2 Tasks Visualization.
We show the samples and detailed annotations of the three tasks in Figure 16.
0.C.1.3 Case Analysis.
Figure 16 presents the results of the Qwen3-VL model on the Spatial Referring task. We observe that existing VLMs perform reasonably well on in-plane spatial referring relations (e.g., “between A and B” as illustrated in the first row). However, when it comes to relations that require understanding positions across different height levels, such as “beneath” or “between A and B” where A and B are situated at different heights, the models fail to grasp the correct spatial information.
0.C.2 Human-centric World Model
0.C.2.1 Gaussian Splatting 4D Reconstruction
@@
During optimization, all input videos are processed at a resolution of 960×720. We represent each scene with 40k dynamic Gaussians and 100k static Gaussians and use B = 10 shared SE(3) motion bases for all points. We optimize the model with Adam, running 1k iterations for the initialization stage and 500 training epochs, with learning rates of for Gaussian means and SE(3) bases, for scales, for rotations, and for colors, opacities, and motion coefficients. Training on a 300-frame sequence takes about 2 hours on a single A100 GPU and supports real-time rendering at around 140 FPS.
0.C.2.2 Addition Application: Language-conditioned Video Generation
This section evaluates the contribution of our proposed WIYH dataset to enhancing video generation performance in human-centric manipulation tasks. We fine-tuned two baseline models, transformer-based CogVideo [hong2022cogvideo] and UNet-based DynamiCrafter [xing2024dynamicrafter], initialized with their original weights, using WIYH under the image-and-text-to-video setting. During fine-tuning, all input videos were resized to 480p resolution. We randomly select 90% of the data as the training set and the remaining 10% as the test set.
Quantitative results are summarized in Table 7. We use four metrics from VBench [huang2024vbench]—consistency, smoothness, dynamic (motion intensity), and quality in order to assess motion plausibility, temporal coherence, and visual fidelity of the video generation models. Following VBench [huang2024vbench], the evaluation metrics are as follows:
After fine-tuning with the WIYH dataset, both baseline models show clear improvements across all metrics, with the most significant gains observed in dynamics and temporal consistency.
| CogVideo | DynamiCrafter | |||
| Metric | w/o WIYH | w/ WIYH | w/o WIYH | w/ WIYH |
| Consistency | 88.2 (+7.6) | 91.5 (+6.2) | ||
| Smoothness | 98.4 (+4.9) | 98.6 (+1.2) | ||
| Dynamic | 78.5 (+15.6) | 89.6 (+14.8) | ||
| Quality | 67.6 (+7.5) | 70.6 (+7.1) | ||
Appendix 0.D Details of Human-centric Manipulation
In this section, we present the details of our co-training experiments in Section 5, Human-centric Manipulation, of the main text. As shown in Figure 18, we demonstrate 11 task variants under two scenarios: Single-Object Scene and Multi-Object Cluttered Scene. Figures 19 and 20 illustrate the success rates and failure causes for each task in these two scenarios, respectively.
As shown in Figures 19 and 20, on most tasks, the incorporation of human data leads to a general increase in success rate. In the Single-Object Scene, the diverse action distribution of human data enhanced the policy’s generalization capability to unseen initial states, enabling successful grasping at different positions or over long distances. In the Multi-Object Cluttered Scene, policies trained solely on robot data demonstrated almost no generalization ability in complex scenarios. When factors such as object height or layout changed, policies trained only on robot data failed, whereas the inclusion of human data gradually enabled success by effectively introducing both the action and observation domains of human data into the robot dataset. Furthermore, the consistently low success rates across all methods in certain tasks highlight important directions for future research: how to further improve the collection precision of human data and how to design efficient algorithms that can better transfer knowledge from human data to dexterous manipulation tasks.
Appendix 0.E Ethics and Privacy Statement